Circular No. 5710

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

4U 0614+09 S. Brandt, N. Lund and A. J. Castro-Tirado, Danish Space Research

Institute, report: “An x-ray burst from the low-mass x-ray binary 4U 0614+09 has been detected by the WATCH all-sky monitor on the EURECA satellite. The burst started on Feb. 17.46513 UT and in eight seconds reached a peak flux of 15 Crab in the 6-15-keV energy band. The flux decayed with an e-folding time of ∼ 15 seconds.”

1993 BW_{2}

Further precise positions have been reported as follows:

1993 UT  α_{2000}  δ_{2000}  Observer
Feb. 2.98084 6^{h}44^{m}00.^{s}54 +39^{o}58′44′′.4 Bendjoya
2.98582 6 44 01.09 +39 59 03.3  _{”}
2.98848 6 44 01.44 +39 59 14.2  _{”}
18.14792 7 16 57.67 +49 45 49.9 McCrosky
18.15744 7 16 58.86 +49 45 59.8  _{”}

P. Bendjoya and A. Maury (Observatoire de la Côte d’Azur). 0.9-m

CERGA Schmidt + CCD. Reduction by Maury and J. B. Edmond. R. E. McCrosky and C.-Y. Shao (Oak Ridge). 1.5-m reflector + CCD.

Improved orbital elements from 12 observations Jan. 28-Feb. 18:

T = 1992 Nov. 20.1158 TT  ω = 287.^{o}2101
e = 0.304419  Ω = 121.2250 }2000.0
q = 0.926126 AU  i = 21.8299

a = 1.331442 AU n^{o} = 0.6415352 P = 1.536 years

1993 TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  V
Feb. 17 7^{h}14.^{m}43 +49^{o}23.^{′}6 0.323 1.218 129.^{o}1 39.^{o}0 17.0
22 7 25.56 +50 43.8 0.369 1.242 125.8 40.2 17.4
27 7 36.87 +51 28.9 0.416 1.266 122.9 41.1 17.7
Mar. 4 7 48.32 +51 48.0 0.465 1.289 120.1 41.7 18.0
9 7 59.84 +51 47.3 0.514 1.312 117.6 42.1 18.3
14 8 11.35 +51 31.1 0.564 1.335 115.1 42.4 18.5
19 8 22.84 +51 02.3 0.614 1.357 112.8 42.5 18.8
24 8 34.28 +50 23.1 0.666 1.379 110.5 42.6 19.0

1993 February 18 Brian G. Marsden